The owners of Speedy Logistics are considering where to locate their new facility. They have narrowed down their search to two locations and have decided to use a factor rating method to make their decision. They have listed the factors they consider important and assigned a factor score to each location based on the five-point scale shown in Table 4.

Table 4

Factor Factor weight location 1 location2
Proximity to airport 40 5 3
Proximity to road access 30 4 1
Proximity to labour source 10 3 5
Size of facility 20 2 4

a. Determine the better location using the factor rating method.

b. A security company wants to locate its headquarters in an urban area in such a way that it minimizes the total distance travelled by its vehicles from the sub branches to the headquarters during any week. The Table 5 shows the coordinates for each branch and the number of trips that will be made per week.

Table 5

NO.OF TRIPS X coordinates Y coordinates
Location 1 3 2 10
Location 2 2 6 8
Location 3 4 4 9
Location 4 1 9 5
Location 5 3 8 1

i. Using the centre of gravity location technique, calculate the coordinates for the location of the headquarters in order to minimize total distance travelled.

c. The management of Sandra Ltd. is considering two different locations for a plant to produce radios. The fixed and variable costs associated with each location are given in the table below.

LOCATION FIXED COST VARIABLE COST
A 600,000 $30/UNIT
B 800,000 $20/UNIT

i. Write an equation for the total costs of each location.

ii. Over which volume of output will each location be preferred?
Answer #1

I have answered only the first question because, the remaining parts are noway related to the first question. Please post them separately and individually. Please make a note that Revised Chegg Guidelines seriously restrict us in answering only one question! Violating the guidelines will get my account suspended.
So, please understand and like this answer. Thank you!

Question 1.  Determine the better location using the factor rating method.

Factor Score at each location

Factor

Factor weight

Location 1

Location 2

Proximity to airport

40

5

3

Proximity to road access

30

4

1

Proximity to labour source

10

3

5

Size of facility

20

2

4

Weighted factor score for each location= sum of (factor weight*factor score for the location) for all 4 factors

Score for location 1 = 40 * 5 + 30 * 4 + 10 * 3 + 20 * 2

= 200 + 120 + 30 + 40

= 390.

Score for location 2 = 40 * 3 + 30 * 1 + 10 * 5 + 20 * 4

= 120 + 30 + 50 + 80

= 280.

Hence, location 1 is better (As it has higher score).